OpenStudy (anonymous):

Factoring by grouping : 48x^y + 20xz + 12xy +5z

OpenStudy (anonymous):

4x^2 + 20x + 5y +xy =4x^2+20x+xy+5y =4x(x+5)+y(x+5) =(4x+y)(x+5) . 48x^2y + 20xz +12xy + 5z 48x^2y+12xy+20xz+5z 12xy(4x+1)+5x(4x+1) (12xy+5x)(4x+1)

OpenStudy (anonymous):

hope that help

OpenStudy (anonymous):

thank you. it did :), im supposed to be doing this online class for this & they do not know how to explain it.

OpenStudy (anonymous):

same go for me but it dont look il be geting much help with my good luck

OpenStudy (anonymous):

what do you need help with ?

OpenStudy (anonymous):

MODELING - Complementary Angles The difference between the measures of two complementary angles is 16°. Determine the measures of the two angles.

OpenStudy (anonymous):

x+y=90 x-y=16 add them: 2x = 106 x = 53 Get y: y = 90-53 = 37 They are 37 degrees and 53 degrees

OpenStudy (anonymous):

thank you